# A Simulation Framework for the LiteBIRD Instruments

### v1

@@ 2 Requirements
The main scientific goal of LiteBIRD is to set an upper limit to the value $\delta r$, the uncertainty on the scalar-to-ratio parameter $r$. Thus, E2E simulations should go through the following steps:
@@
This list prompts us to make two crucial remarks. First, the data acquisition process is modeled only in the first two steps of the procedure: technically speaking, steps 3–6 are data reduction tasks, which are the same for simulated and real data. However, to ensure that the results of E2E simulations can be accurately interpreted and fed into the next stages of the analysis, we had to include a few data-reduction modules (map-makers) in our simulation framework, as explained in Section 4.2. The second remark is that $N$ Monte Carlo realizations are needed to estimate error bars for the scientific parameters produced by E2E simulations. This increases the processing power required to run the simulations, necessitating the framework to be optimized for both speed and memory.
@@ 3.2.1 Scientific samples
Figure 1: The way LBS can split a 2D matrix containing a TOD, as well as any other attribute associated with each detector, like the noise level, the FWHM of the beam, etc. The first row illustrates how splits are made when the code is run in serial mode, i.e., without MPI. The parameters n_blocks_det and n_blocks_time specify the number of splits between the detectors and along the time axis, respectively. The bottom row shows how TODs are split when there are four MPI processes. Attributes can be copied over many MPI processes if these processes handle the same detectors. (The latter is the case of the two panels on the right.)
@@ 3.2.2 Pointing information and Half-Wave Plate angles
Beyond TODs, the collection of pointing information also requires consistent memory allocations. Pointing information specifies the direction and orientation of each detector’s beam axis at the time when each sample was measured, and it is typically expressed through three angles: colatitude, longitude, and orientation (tilt). Consequently, the pointing matrix for $N$ detectors and $M$ time samples is a $N\times M\times 3$ array. For MPI applications, LBS stores [^9] As it is the case for CMB space surveys, the nominal scanning strategy is a composition of rotations with constant angular speed. In simulation codes, the typical approach is to encode these rotations as lowly-sampled quaternions, e.g., one per second or even less, depending on the angular speeds, and then use slerp operations to compute pointing angles at the same sampling frequency as scientific data. Since version 0.13.0, LBS can compute pointings for each detector on the fly from quaternions, thus avoiding the need to store the pointing matrices and significantly reducing memory usage. these matrices using the same distribution strategy as the TOD matrices.
@@ 3.2.3 I/O
One of the purposes of LBS is to produce simulated data to be fed into data reduction pipelines, thereby validating the latter. To enable this interaction, LBS provides tools that save the timelines and the pointing information in HDF5 files. In version 0.11.0, LBS ensures that every MPI process saves its TODs and pointings in separate files; thus, a program running on $N$ MPI processes will save $N$ files. To save disk space, the user can use a flag to instruct LBS to store the low-sampled quaternions used for computing the pointings instead of the full $N\times M\times 3$ pointing matrix.
@@ 4.1.2 Scanning strategy
  • The Sun-Earth axis rotates around the Sun with a period of 365365~{}365365 days.
@@ 4.1.4 Solar dipole
LBS implements several simplifications of the formula that we call “dipole models”; they are reported [^19] https://litebird-sim.readthedocs.io/en/latest/dipole.html in the User’s manual.
@@ 5.3 Reproducibility
Finally, there is the possibility that the code has some state that is not preserved across separate runs. There is no reliable solution that can work 100% of the time to prevent hidden state [^24] In principle, using functional languages like Haskell or Clojure and purely functional data structures might prevent this kind of error. However, no language is perfectly functional and thus the problem would still be unsolved. from altering the results of a computation. However, a common source of confusion is the seed used to generate pseudo-random numbers, as it is often initialized to a value derived from the system clock. LBS requires that the parameter random_seed accepted by the constructor of the class Simulation be always provided. The seed can be None or an integer number; setting it to None uses the internal clock and thus prevents the reproducibility of the results; however, we fell the fact that random_seed=None must be spelled explicitly in the source code to enable this behaviour is a good-enough “red flag”. Instead, by setting random_seed to an int, the results of separate runs of a script will produce the same sequence of pseudo-random numbers. The seed can be changed by using sim.init_random, i.e., the same function used at the end of the Simulation constructor to set up the RNG:
@@ 6.1 Setting up the simulation
The center point of LBS is the Simulation class, which should be instantiated in any pipeline built using this framework. The class serves as a container for several analysis modules available to the user, and it tracks both the inputs provided by the user and the output data generated by the simulation itself. Several information about the simulation need to be shared among different modules: for instance, the launch date and the duration of the simulation impact both the code that allocates the TOD, because it needs to know how many samples to allocate in memory, and the code that simulates the presence of moving sources in the sky, because it requires to compute the ephemerides of the planets.
@@ 6.5 Scanning strategy
Each pointing is encoded as a set of three angles: the colatitude and longitude in Ecliptic coordinates (in radians), and the orientation angle of the detector projected in the same coordinate system; the latter is used to determine the amount of polarized light coming from the sky that is measured by the detector. It is possible to compute other parameters related to the spacecraft’s motion, but we will discuss them in Section 6.8, where we will describe how LBS simulates the CMB dipole signal.
@@ 6.10 Map-making
[2025-04-11 20:30:09,554 INFO MPI#$0$] Destriper CG iteration 1/100,
@@
[2025-04-11 20:30:09,714 INFO MPI#$0$] Destriper CG iteration 2/100,

### v2

@@ 2 Requirements
The main scientific goal of LiteBIRD is to set an upper limit to the value $\delta r$, the uncertainty on the scalar-to-ratio parameter $r$, under the assumption of a fiducial model with $r=0$ [1]. Thus, E2E simulations should go through the following steps:
@@
This list prompts us to make two crucial remarks. First, the data acquisition process is modeled only in the first two steps of the procedure: technically speaking, steps 3–6 are data reduction tasks, which are the same for simulated and real data. However, to ensure that the results of E2E simulations can be accurately interpreted and fed into the next stages of the analysis, we had to include a few data-reduction modules (map-makers) in our simulation framework, as explained in Section 4.2. The second remark is that a certain number $N$ of Monte Carlo realizations is required to estimate error bars for the scientific parameters produced by E2E simulations. This increases the processing power required to run the simulations, necessitating the framework to be optimized for both speed and memory.
@@ 3.2.1 Scientific samples
Figure 1: The way LBS can split a 2D matrix containing a TOD, as well as any other attribute associated with each detector, like the noise level, the FWHM of the beam, etc. The first row illustrates how splits are made when the code is run in serial mode, i.e., without MPI. The parameters n_blocks_det and n_blocks_time specify the number of splits between the detectors and along the time axis, respectively. The bottom right panel shows how TODs are split when there are four MPI processes. Attributes can be copied over many MPI processes if these processes handle the same detectors. (The latter is the case of the two panels on the right.)
@@ 3.2.2 Pointing information and Half-Wave Plate angles
Beyond TODs, the collection of pointing information also requires consistent memory allocations. Pointing information specifies the direction and orientation of each detector’s beam axis at the time when each sample was measured, and it is typically expressed through three angles: colatitude, longitude, and orientation (tilt). Consequently, the pointing matrix for $N$ detectors and $M$ time samples is a $N\times M\times 3$ array. For MPI applications, LBS stores [^9] As it is the case for CMB space surveys, the nominal scanning strategy is a composition of rotations with constant angular speed. In simulation codes, the typical approach is to encode these rotations as lowly-sampled quaternions, e.g., one per second or even less, depending on the angular speeds, and then use spherical linear operations (“slerp”) to compute pointing angles at the same sampling frequency as scientific data. Since version 0.13.0, LBS can compute pointings for each detector on the fly from quaternions, thus avoiding the need to store the pointing matrices and significantly reducing memory usage. these matrices using the same distribution strategy as the TOD matrices.
@@ 3.2.3 I/O
One of the purposes of LBS is to produce simulated data to be fed into data reduction pipelines, thereby validating the latter. To enable this interaction, LBS provides tools that save the timelines and the pointing information in HDF5 files. In version 0.11.0, LBS ensures that every MPI process saves its TODs and pointings in separate files; thus, a program running on $N$ MPI processes will save $2N$ files. To save disk space, the user can use a flag to instruct LBS to store the low-sampled quaternions used for computing the pointings instead of the full $N\times M\times 3$ pointing matrix.
@@ 3.3 Provenance tracking
A simplified view of the public PTEP IMo bundled with LBS 0.11.0 is illustrated in Fig. 2. The relevant information is stored in the so-called “quantities”, which are represented as oval nodes with a gray background; all the quantities shown in the figure are JSON records containing several parameters that describe the “entity” to which they refer:
  • A scanning_parameters quantity contains the parameters of the nominal scanning strategy: spinning speed, angles of the rotation axes, etc.
  • A instrument_info quantity contains general information about the instrument, such as the number of frequency channels or the angle between the boresight of the focal plane and the spin axis of the spacecraft.
  • A detector_info quantity contains several synthetic parameters that describe the performance of a single detector, such as the NET and $1/f$ noise characteristics.
  • A channel_info quantity contains the average values of several quantities associated with detectors operating at the same nominal frequency. They are the same as those found in the detector_info objects, and are used for rough simulations where the detailed behavior of every single detector is not relevant.
As explained before, information is retrieved from the IMo through paths, which are plain Python strings. Thus, to retrieve the JSON dictionary associated with detector 000_000_003_QA_040_T, the path to be passed to LBS is
satellite/LFT/L1-040/000_000_003_QA_040_T
Note that, as LBS refers to objects through their full path, it can handle any topology of the IMo. The official IMo database, which is not publicly available, uses a more complicated layout.
Figure 2: Simplified view of the PTEP IMo bundled with LBS 0.11.0. Boxes represent “entities”, i.e., the nodes that provide the structure (topology) of the tree, while gray ovals represent the “quantities” that contain actual data. (A third level is represented by “data files”, which are different versions of a quantity and are similar to commits in a Git repository.) The image only shows a few entities and quantities of the whole tree: for each instrument (LFT, MFT, HFT), only two channels are represented, and each channel shows only two detectors.
@@ 4.1.2 Scanning strategy
  • The Sun-Earth axis rotates around the Sun with a period of 365~365 days.
@@ 4.1.4 Solar dipole
LBS implements several simplifications of Eq. (4.2) that we call “dipole models”; they are reported [^19] https://litebird-sim.readthedocs.io/en/latest/dipole.html in the User’s manual.
@@ 5.3 Reproducibility
Finally, there is the possibility that the code has some state that is not preserved across separate runs. There is no reliable solution that can work 100% of the time to prevent hidden state [^24] In principle, using functional languages like Haskell or Clojure and purely functional data structures might prevent this kind of error. However, no language is perfectly functional and thus the problem would still be unsolved. from altering the results of a computation. However, a common source of confusion is the seed used to generate pseudo-random numbers, as it is often initialized to a value derived from the system clock. LBS requires that the parameter random_seed accepted by the constructor of the class Simulation be always provided. The seed can be None or an integer number; setting it to None uses the internal clock and thus prevents the reproducibility of the results; however, we feel the fact that random_seed=None must be spelled explicitly in the source code to enable this behaviour is a good-enough “red flag”. Instead, by setting random_seed to an int, the results of separate runs of a script will produce the same sequence of pseudo-random numbers. The seed can be changed by using sim.init_random, i.e., the same function used at the end of the Simulation constructor to set up the RNG:
@@ 6.1 Setting up the simulation
The center point of LBS is the Simulation class, which should be instantiated in any pipeline built using this framework. The class serves as a container for several analysis modules available to the user, and it tracks both the inputs provided by the user and the output data generated by the simulation itself. Several pieces of information about the simulations need to be shared among different modules: for instance, the launch date and the duration of the simulation impact both the code that allocates the TOD, because it needs to know how many samples to allocate in memory, and the code that simulates the presence of moving sources in the sky, because it is required to compute the ephemerides of the planets.
@@ 6.5 Scanning strategy
Each pointing is encoded as a set of three angles: the colatitude and longitude in Ecliptic coordinates (in radians), and the orientation angle of the detector projected in the same coordinate system; the latter is used to determine which linear polarization component is measured by the detector, thereby enabling the reconstruction of the Stokes $Q$ and $U$ parameters. It is possible to compute other parameters related to the spacecraft’s motion, but we will discuss them in Section 6.8, where we will describe how LBS simulates the CMB dipole signal.
@@ 6.10 Map-making
[2025-04-11 20:30:09,554 INFO MPI#0000] Destriper CG iteration 1/100,
@@
[2025-04-11 20:30:09,714 INFO MPI#0000] Destriper CG iteration 2/100,